Find the sum: (-8/19) + (-2/57) Find the sum: (-8/19) + (-2/57)

Find the sum:     (-8/19) + (-2/57)
Find the sum:     (-8/19) + (-2/57)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have
= -8/19–2/57
Take the LCM of the denominators of the given rational numbers.
LCM of 19 and 57 is 57
Express each of the given rational numbers with the above LCM as the common denominator.
Now,
(-8/19)= [(-8×3)/ (19×3)] = (-24/57)
(-2/57)= [(-2×1)/ (57×1)] = (-2/57)
Then,
= (-24/57) - (2/57)
= (-24 - 2)/57
= (-26/57) ... [∵denominator is same in both the rational numbers]

Thanks